### v2

# Lê modules and hypersurfaces with one-dimensional singular sets

Abstract.
By using our previous results on Lê modules and an upper-bound on the betti numbers which we proved with Lê, we investigate the cohomology of Milnor fibers and the internal local systems given by the vanishing cycles of hypersurfaces with one-dimensional singular sets.
2020 Mathematics Subject Classification:
32S25, 32S15, 32S55
1. Introduction
Let $\mathcal{U}$} be a connected open neighborhood of the origin in ${\mathbb{C}}^{n+1}$, let ${\mathbf{z}}:=(z_{0},\dots,z_{n})$ be coordinates on $\mathcal{U}$}, and let $f:({\mathcal{U}},{\mathbf{0}})\rightarrow({\mathbb{C}},0)$ be a non-constant, reduced complex analytic function. After possibly choosing $\mathcal{U}$} smaller, the critical locus, $\Sigma f$, of $f$ is equal to the singular set $\Sigma V(f)$ of the hypersurface $V(f):=f^{-1}(0)$. In this paper, we assume that $\dim_{\mathbf{0}}\Sigma f=1$ and assume, possibly after an analytic change of coordinates, that $z_{0}$ is generic enough so that $\dim_{\mathbf{0}}\Sigma(f_{|_{V(z_{0})}})=0$ and so that, for all irreducible components $C$ of $\Sigma f$ at $\mathbf{0}$}, $\left(C\cdot V(z_{0})\right)_{\mathbf{0}}={\operatorname{mult}}_{\mathbf{0}}C$.
In many places, we have discussed the Lê numbers of hypersurface singularities; see, for instance, [8], [9], [10], [7], and [11]. In our current setting, there are only two possibly non-zero Lê numbers at the origin, $\lambda^{0}_{f,{\mathbf{z}}}({\mathbf{0}})$ and $\lambda^{1}_{f,{\mathbf{z}}}({\mathbf{0}})$, which we will henceforth denote simply as $\lambda^{0}$ and $\lambda^{1}$; these are guaranteed to exist, since our genericity assumption on $z_{0}$ implies the coordinates $\mathbf{z}$} are prepolar for $f$ at $\mathbf{0}$}. In the next section, we will recall the definitions of $\lambda^{0}$ and $\lambda^{1}$, but here let us just say that $\lambda^{0}$ and $\lambda^{1}$ are effectively calculable, by hand in many examples and by computer in harder examples.
[…]
There will be subcases in which we also need to consider
[…]
where the vertical lines indicate that we are considering $\Sigma f$ with its reduced structure, i.e., as a set.
We give our results here in the introduction, but will defer the proofs until Section 3.
As we are assuming that $\dim_{\mathbf{0}}\Sigma f=1$, $\lambda^{1}\neq 0$, so let us start with the case:
Proposition 1.1.
Suppose that $\lambda^{0}=0$.
Then, at $\mathbf{0}$}, $\Sigma f$ is smooth, transversely intersected by $V(z_{0})$, and $f_{z_{0}}(z_{1},\dots,z_{n})=f(z_{0},\dots,z_{n})$ defines a 1-parameter family of isolated singularities along $C$ with constant Milnor number $\mu^{\circ}$.
Furthermore,
$\widetilde{H}^{n-1}(F_{f,\mathbf{0}};\ {\mathbb{Z}})\cong{\mathbb{Z}}^{\mu^{\circ}}\qquad\textnormal{ and }\qquad\widetilde{H}^{n}(F_{f,\mathbf{0}};\ {\mathbb{Z}})=0.$
Proposition 1.2.
Suppose $\lambda^{1}=1$ and $\lambda^{0}\geq 1$.
Then,
$\widetilde{H}^{n-1}(F_{f,\mathbf{0}};\ {\mathbb{Z}})=0\qquad\textnormal{ and }\qquad\widetilde{H}^{n}(F_{f,\mathbf{0}};\ {\mathbb{Z}})\cong{\mathbb{Z}}^{\lambda^{0}-1},$
and $\lambda^{0}\geq 2$.
Proposition 1.3.
Suppose $\lambda^{0}=1$ and $\lambda^{1}\geq 2$.
Then
$\widetilde{H}^{n-1}(F_{f,\mathbf{0}};\ {\mathbb{Z}})\cong{\mathbb{Z}}^{\lambda^{1}-1}\qquad\textnormal{ and }\qquad\widetilde{H}^{n}(F_{f,\mathbf{0}};\ {\mathbb{Z}})=0.$
Furthermore, ${\operatorname{mult}}_{\mathbf{0}}f=2$ and $m=2$, and if $r=1$, then $\lambda^{1}=2$, and so $\widetilde{H}^{n-1}(F_{f,\mathbf{0}};\ {\mathbb{Z}})\cong{\mathbb{Z}}$.
Proposition 1.4.
Suppose $\lambda^{1}=2$ and $\lambda^{0}\geq 2$.
Then, in fact, $\lambda^{0}\geq 3$ and there are two cases:
Case 1:
$\widetilde{H}^{n-1}(F_{f,\mathbf{0}};\ {\mathbb{Z}})=0\qquad\textnormal{ and }\qquad\widetilde{H}^{n}(F_{f,\mathbf{0}};\ {\mathbb{Z}})\cong{\mathbb{Z}}^{\lambda^{0}-2}\oplus\big({\mathbb{Z}}/b{\mathbb{Z}}\big)$
for some $b\geq 1$.
Case 2:
$\widetilde{H}^{n-1}(F_{f,\mathbf{0}};\ {\mathbb{Z}})\cong{\mathbb{Z}}\qquad\textnormal{ and }\qquad\widetilde{H}^{n}(F_{f,\mathbf{0}};\ {\mathbb{Z}})\cong{\mathbb{Z}}^{\lambda^{0}-1}.$
Furthermore, if we also assume that $m=1$, then we are in Case (1) above and
$\widetilde{H}^{n-1}(F_{f,\mathbf{0}};\ {\mathbb{Z}})=0\qquad\textnormal{ and }\qquad\widetilde{H}^{n}(F_{f,\mathbf{0}};\ {\mathbb{Z}})\cong{\mathbb{Z}}^{\lambda^{0}-2}\oplus\big({\mathbb{Z}}/b{\mathbb{Z}}\big),$
where either $b=1$ (so there is no torsion) or $b$ is a product of prime powers where the primes are either $3$ or congruent to $1$ modulo $6$.
Proposition 1.5.
Suppose $\lambda^{0}=2$ and $\lambda^{1}\geq 3$.
There are two cases:
$\widetilde{H}^{n-1}(F_{f,\mathbf{0}};\ {\mathbb{Z}})\cong{\mathbb{Z}}^{\lambda^{1}-2}\qquad\textnormal{ and }\qquad\widetilde{H}^{n}(F_{f,\mathbf{0}};\ {\mathbb{Z}})\cong{\mathbb{Z}}/b{\mathbb{Z}}$
for some $b\geq 1$.
$\widetilde{H}^{n-1}(F_{f,\mathbf{0}};\ {\mathbb{Z}})\cong{\mathbb{Z}}^{\lambda^{1}-1}\qquad\textnormal{ and }\qquad\widetilde{H}^{n}(F_{f,\mathbf{0}};\ {\mathbb{Z}})\cong{\mathbb{Z}}.$
Furthermore, if $m=2$, then we are in Case (1) above and either $b=1$ or $b$ is a product of prime powers where the primes are either $3$ or congruent to $1$ modulo $6$.
Proposition 1.6.
Suppose $\lambda^{0}=\lambda^{1}=3$.
Then
$\widetilde{H}^{n-1}(F_{f,\mathbf{0}};\ {\mathbb{Z}})\cong{\mathbb{Z}}^{2}\qquad\textnormal{ and }\qquad\widetilde{H}^{n}(F_{f,\mathbf{0}};\ {\mathbb{Z}})\cong{\mathbb{Z}}^{2}.$
Furthermore, if $r=1$, then $m=1$, i.e., $\Sigma f$ is smooth and transversely intersected by $V(z_{0})$ at $\mathbf{0}$}
This paper is organized as follows. In Section 2, we will give definitions and prior results. Section 3 is devoted to proving the propositions/cases given above. Section 4 contains examples and mentions when we have none. Section 5 contains concluding remarks.
2. Definitions and prior results
We should be clear from the beginning that, when we look at intersection numbers, we deal with analytic cycles, not cycle classes, and our intersection theory is the simple theory using proper intersections inside the complex manifold $\mathcal{U}$}; one can find this in [5] in Section 1.5, Section 8.2, and Example 11.4.4.
Our assumption that
$\dim_{\mathbf{0}}\Sigma\big(f_{|_{V(z_{0})}}\big)\ =\ \dim_{\mathbf{0}}V\left(z_{0},\frac{\partial f}{\partial z_{1}},\dots,\frac{\partial f}{\partial z_{n}}\right)\ =\ 0$
is precisely equivalent to saying that $V\left(\frac{\partial f}{\partial z_{1}},\dots,\frac{\partial f}{\partial z_{n}}\right)$ is purely $1$-dimensional (and not empty) at the origin and is properly intersected at the origin by $V(z_{0})$.
In terms of analytic cycles,
$V\left(\frac{\partial f}{\partial z_{1}},\dots,\frac{\partial f}{\partial z_{n}}\right)\ =\ \Gamma^{1}_{f,{\mathbf{z}}}\ +\ \Lambda^{1}_{f,{\mathbf{z}}},$
where $\Gamma^{1}_{f,{\mathbf{z}}}$ is the (possibly non-reduced) relative polar curve of $f$ of Hamm, Lê, and Teissier, which consists of components not contained in $\Sigma f$, and $\Lambda^{1}_{f,{\mathbf{z}}}$ is the $1$-dimensional Lê cycle, which consists of components which are contained in $\Sigma f$. See Definition 1.11 of [10]. Note that $V\left(\frac{\partial f}{\partial z_{0}}\right)$ necessarily intersects $\Gamma^{1}_{f,{\mathbf{z}}}$ properly at $\mathbf{0}$}, and that $V(z_{0})$ intersects $\Lambda^{1}_{f,{\mathbf{z}}}$ properly at $\mathbf{0}$} by our assumption.
Letting $C$’s denote the underlying reduced components of $\Sigma f$ at $\mathbf{0}$}, at the origin, we have
$\Lambda^{1}_{f,{\mathbf{z}}}\ =\ \sum_{C}\mu^{\circ}_{C}\,C,$
where $\mu^{\circ}_{C}$ is the Milnor number of $f$ restricted to a generic hyperplane slice, at a point $\mathbf{p}$} on $C-\{{\mathbf{0}}\}$ close to $\mathbf{0}$}. See Remark 1.19 of [10].
The intersection numbers $\left(\Gamma^{1}_{f,{\mathbf{z}}}\cdot V\left(\frac{\partial f}{\partial z_{0}}\right)\right)_{\mathbf{0}}$ and $\left(\Lambda^{1}_{f,{\mathbf{z}}}\cdot V(z_{0})\right)_{\mathbf{0}}$ are, respectively, the Lê numbers $\lambda^{0}=\lambda^{0}_{f,{\mathbf{z}}}$ and $\lambda^{1}=\lambda^{1}_{f,{\mathbf{z}}}$ (at the origin). See Definition 1.11 of [10]. Note that
$\lambda^{1}=\sum_{C}\left(C\cdot V(z_{0})\right)_{\mathbf{0}}\mu^{\circ}_{C}.$
We wish now to discuss the main result of [7]. Note that $\widetilde{H}^{n-1}(F_{f,{\mathbf{0}}};{\mathbb{Z}})$ is free abelian, while $\widetilde{H}^{n}(F_{f,{\mathbf{0}}};{\mathbb{Z}})$ may have torsion. We write
$\widetilde{H}^{n-1}(F_{f,{\mathbf{0}}};{\mathbb{Z}})\cong{\mathbb{Z}}^{\tilde{b}_{n-1}}\ \textnormal{ and }\ \widetilde{H}^{n}(F_{f,{\mathbf{0}}};{\mathbb{Z}})\cong{\mathbb{Z}}^{\tilde{b}_{n}}\oplus T,$
where $T$ consists of pure torsion. For each prime number $\mathfrak{p}$, we write $\tau_{\mathfrak{p}}$ for the number of cyclic summands of $T$ with $\mathfrak{p}$-torsion, i.e., the number of direct summands of $T$ of the form ${\mathbb{Z}}/\mathfrak{p}^{k}{\mathbb{Z}}$.
The Universal Coefficient Theorem tells us that
$\widetilde{H}^{n-1}(F_{f,{\mathbf{0}}};{\mathbb{Z}}/\mathfrak{p}{\mathbb{Z}})\cong\left({\mathbb{Z}}/\mathfrak{p}{\mathbb{Z}}\right)^{\tilde{b}_{n-1}+\,\tau_{\mathfrak{p}}}\ \textnormal{ and }\ \widetilde{H}^{n}(F_{f,{\mathbf{0}}};{\mathbb{Z}}/\mathfrak{p}{\mathbb{Z}})\cong\left({\mathbb{Z}}/\mathfrak{p}{\mathbb{Z}}\right)^{\tilde{b}_{n}+\,\tau_{\mathfrak{p}}}.$
Now Theorem 2.5 of [7] tells us:
Theorem 2.1.
stuff
Either
$\bullet$ $\lambda^{0}=0$, $\Sigma f$ is smooth at $\mathbf{0}$} and $f$ defines a 1-parameter family of isolated singularities with constant Milnor number $\mu$, in which case $\lambda^{0}=0$, $\widetilde{H}^{n-1}(F_{f,\mathbf{0}};\ {\mathbb{Z}})\cong{\mathbb{Z}}^{\mu}$, and $\widetilde{H}^{n}(F_{f,\mathbf{0}};\ {\mathbb{Z}})=0$,
or
$\bullet$ $\lambda^{0}\neq 0$ and, for all primes $\mathfrak{p}$, there is a strict inequality $\tilde{b}_{n-1}+\,\tau_{\mathfrak{p}}<\lambda^{1}$ and, equivalently, $\tilde{b}_{n}+\,\tau_{\mathfrak{p}}<\lambda^{0}$.
We will now define the Lê modules. We fix a base ring $R$ which is a principal ideal domain (e.g., $\mathbb{Z}$} or any field), and consider the shifted constant sheaf (complex) $R_{{\mathcal{U}}}^{\bullet}[n+1]$; this is a perverse sheaf (see, for instance, [2], [6], [4]). It follows from the general theory that the shifted vanishing cycle complex along $f$, $\mathbf{P}^{\bullet}:=\phi_{f}[-1]R_{{\mathcal{U}}}^{\bullet}[n+1]$, is also a perverse sheaf, as are $\psi_{z_{0}}[-1]\mathbf{P}^{\bullet}$ and $\phi_{z_{0}}[-1]\mathbf{P}^{\bullet}$, the shifted nearby and vanishing cycles of $\mathbf{P}^{\bullet}$ along $z_{0}$. Furthermore the origin is an isolated point in the support of $\psi_{z_{0}}[-1]\mathbf{P}^{\bullet}$ and $\phi_{z_{0}}[-1]\mathbf{P}^{\bullet}$, and there is a canonical morphism
$\psi_{z_{0}}[-1]\mathbf{P}^{\bullet}\xrightarrow{{\ \operatorname{can}}_{z_{0}}\ }\phi_{z_{0}}[-1]\mathbf{P}^{\bullet}.$
As discussed in [11], the stalk cohomologies in degree 0 of $\psi_{z_{0}}[-1]\mathbf{P}^{\bullet}$ and $\phi_{z_{0}}[-1]\mathbf{P}^{\bullet}$ (the Lê modules) are free modules of ranks $\lambda^{1}$ and $\lambda^{0}$, respectively, i.e.,
$H^{0}(\psi_{z_{0}}[-1]\mathbf{P}^{\bullet})_{\mathbf{0}}\cong R^{\lambda^{1}}\qquad\textnormal{ and }\qquad H^{0}(\phi_{z_{0}}[-1]\mathbf{P}^{\bullet})_{\mathbf{0}}\cong R^{\lambda^{0}}$
and the map ${\operatorname{can}}_{z_{0}}$ induces the map $\partial$ from the introduction, i.e., an $R$-module homomorphism $\partial:R^{\lambda^{1}}\rightarrow R^{\lambda^{0}}$ such that
$\ker\partial\cong\widetilde{H}^{n-1}(F_{f,\mathbf{0}};\ R)\qquad\textnormal{ and }\qquad{\operatorname{coker}}\,\partial\cong\widetilde{H}^{n}(F_{f,\mathbf{0}};\ R).$
In the context of a singular set of dimension 1, and recalling that $m:={\operatorname{mult}}_{\mathbf{0}}|\Sigma f|$, what we prove in [11] is that
Theorem 2.2.
stuff
  • There exists a $\mathbb{Z}$}-module homomorphism $\partial:{\mathbb{Z}}^{\lambda^{1}}\rightarrow{\mathbb{Z}}^{\lambda^{0}}$ such that
    $\ker\partial\cong\widetilde{H}^{n-1}(F_{f,\mathbf{0}};\ {\mathbb{Z}})\qquad\textnormal{ and }{\operatorname{coker}}\,\partial\cong\widetilde{H}^{n}(F_{f,\mathbf{0}};\ {\mathbb{Z}}),$
    where $\widetilde{H}^{\bullet}$ denotes reduced cohomology.
  • For $j=0,1$, the Milnor monodromy of $f$ induces a Lê-Milnor monodromy automorphism $\alpha_{j}$ from ${\mathbb{Z}}^{\lambda^{j}}$ to itself which is compatible with the map $\partial$, i.e., $\partial\circ\alpha_{1}=\alpha_{0}\circ\partial$.
  • All of the eigenvalues of each $\alpha_{j}$ are roots of unity, and so the characteristic polynomial of each $\alpha_{j}$ factors into cyclotomic polynomials.
  • The traces, $\operatorname{tr}\alpha_{j}$, of $\alpha_{0}$ and $\alpha_{1}$ are given by
    $\operatorname{tr}\alpha_{0}=(-1)^{n}(m-1)\qquad\textnormal{ and }\qquad\operatorname{tr}\alpha_{1}=(-1)^{n}\,m.$
    As all of the eigenvalues of $\alpha_{0}$ are roots of unity, it follows that $m-1\leq\lambda^{0}$, with equality holding if and only if the characteristic polynomial of $\alpha_{0}$ is either $(t-1)^{\lambda^{0}}$ or $(t+1)^{\lambda^{0}}$.
  • if $\mathfrak{p}$ is a prime number, then tensoring $\partial$ with ${\mathbb{Z}}/\mathfrak{p}{\mathbb{Z}}$ yields a linear transformation of ${\mathbb{Z}}/\mathfrak{p}{\mathbb{Z}}$-vector spaces $\partial^{\mathfrak{p}}:\left({\mathbb{Z}}/\mathfrak{p}{\mathbb{Z}}\right)^{\lambda^{1}}\rightarrow\left({\mathbb{Z}}/\mathfrak{p}{\mathbb{Z}}\right)^{\lambda^{0}}$ such that
    $\ker\partial^{\mathfrak{p}}\cong\widetilde{H}^{n-1}(F_{f,\mathbf{0}};\ {\mathbb{Z}}/\mathfrak{p}{\mathbb{Z}})\qquad\textnormal{ and }{\operatorname{coker}}\,\partial^{{}\,\mathfrak{p}}\cong\widetilde{H}^{n}(F_{f,\mathbf{0}};\ {\mathbb{Z}}/\mathfrak{p}{\mathbb{Z}}).$
The perversity of $\mathbf{P}^{\bullet}$ implies, among other things, that, for $\mathcal{U}$} small enough, for all components $C$ of $\Sigma f$, the cohomology sheaf ${\mathbf{H}}^{-1}\big(\mathbf{P}^{\bullet}_{|_{C-\{{\mathbf{0}}\}}}\big)$ is a local system on $C-\{{\mathbf{0}}\}$ with stalk isomorphic to ${\mathbb{Z}}^{\mu^{\circ}_{C}}$.
In classical terms, this is well-known; it says that, for $p\in C-\{{\mathbf{0}}\}$, (i) the reduced cohomology $\widetilde{H}^{*}(F_{f,p})$ is possibly non-zero only in degree n1n-1, (ii) the isomorphism-type of the module $\widetilde{H}^{n-1}(F_{f,p})$ is independent of $p\in C-\{{\mathbf{0}}\}$ and is free abelian of rank that we denoted above by $\mu^{\circ}_{C}$, and (iii) there is a monodromy action/homomorphism $\rho_{C}:\pi_{1}(C-\{{\mathbf{0}}\})\cong{\mathbb{Z}}\rightarrow\operatorname{Aut}({\mathbb{Z}}^{\mu^{\circ}})$. Let $h_{C}$ denote the automorphism $\rho_{C}(1)$.
This internal local system monodromy is sometimes referred to as the vertical monodromy; this is to distinguish it from the Milnor monodromy induced by letting the value of $f$ travel around a small circle.
Restricting our setting to a small neighborhood of the origin, the cosupport condition for perverse sheaves (see also Siersma [13]) implies:
Proposition 2.3.
$H^{-1}(\mathbf{P}^{\bullet})_{\mathbf{0}}\cong\widetilde{H}^{n-1}(F_{f,{\mathbf{0}}};{\mathbb{Z}})$ injects into $\bigoplus_{C}\ker\{\operatorname{id}-h_{C}\}$. In particular, $\tilde{b}_{n-1}\leq\sum_{C}\mu^{\circ}_{C}$.
We have one other result to mention here.
Proposition 2.4.
Suppose that $\tilde{b}_{n-1}$ equals 0 or 1. Then $\lambda^{0}\neq\lambda^{1}$.
Proof.
Suppose that $\tilde{b}_{n-1}=0$ and $\lambda^{0}=\lambda^{1}$. Then we would also have $\tilde{b}_{n}=0$, which would contradict A’Campo’s Theorem in [1] that the Lefschetz number of the monodromy on the Milnor fiber is 0.
Suppose that $\tilde{b}_{n-1}=1$ and $\lambda^{0}=\lambda^{1}$. Then we would also have $\tilde{b}_{n}=1$, which would imply that the traces of the Milnor automorphisms on $\widetilde{H}^{n-1}(F_{f,{\mathbf{0}}};{\mathbb{Q}})$ and $\widetilde{H}^{n}(F_{f,{\mathbf{0}}};{\mathbb{Q}})$ are each $\pm 1$. This would again contradict A’Campo’s Theorem. ∎
Now we are prepared to prove all of the propositions/cases from the introduction.
3. Proofs
Proof of Proposition 1.1. Suppose that $\lambda^{0}=0$.
$\lambda^{0}=0$ if and only if the relative polar curve near the origin is empty as a set, or is 0 as a cycle; it is well-known (see, for instance, Theorem 5.1 of [7]) that this is equivalent to:
$\Sigma f$ is smooth at $\mathbf{0}$}, and so $\Lambda^{1}=\mu^{\circ}C$, for some positive integer $\mu^{\circ}$ and some irreducible curve $C$ which is smooth at $\mathbf{0}$}, $V(z_{0})$ transversely intersects $C$ at $\mathbf{0}$}, and $f_{z_{0}}(z_{1},\dots,z_{n})=f(z_{0},\dots,z_{n})$ defines a 1-parameter family of isolated singularities along $C$ with constant Milnor number $\mu^{\circ}$.
Moreover, these equivalent conditions imply that $\widetilde{H}^{n-1}(F_{f,\mathbf{0}};\ {\mathbb{Z}})\cong{\mathbb{Z}}^{\mu^{\circ}}$ and $\widetilde{H}^{n}(F_{f,\mathbf{0}};\ {\mathbb{Z}})=0$.
Proof of Proposition 1.2. Suppose $\lambda^{1}=1$ and $\lambda^{0}\geq 1$.
In this case, $\Lambda^{1}_{f,{\mathbf{z}}}=C$, where $C$ is smooth at $\mathbf{0}$}, $\mu^{\circ}_{C}=1$, and $V(z_{0})$ transversely intersects $C$ at $\mathbf{0}$}. After an analytic change of coordinates, we may assume that $C$ is a complex line, and we are in the case that Siersma refers to as an isolated line singularity in [12]. As we are assuming that $\lambda^{0}\neq 0$, Siersma proves in this case that $F_{f,{\mathbf{0}}}$ has the homotopy-type of a bouquet of $n$-spheres; the number of $n$-spheres in the bouquet is necessarily $\lambda^{0}-1$.
Therefore,
$\widetilde{H}^{n-1}(F_{f,\mathbf{0}};\ {\mathbb{Z}})=0\qquad\textnormal{ and }\qquad\widetilde{H}^{n}(F_{f,\mathbf{0}};\ {\mathbb{Z}})\cong{\mathbb{Z}}^{\lambda^{0}-1}.$
Note that Proposition 1.6 implies $\lambda^{0}\neq 1$.
In terms of the results discussed in the prior section, the above result on the cohomology of $F_{f,{\mathbf{0}}}$ follows at once from Theorem 2.1.∎
Proof of Proposition 1.3. Suppose $\lambda^{0}=1$ and $\lambda^{1}\geq 2$.
In this case, Theorem 2.1 immediately yields
$\widetilde{H}^{n-1}(F_{f,\mathbf{0}};\ {\mathbb{Z}})\cong{\mathbb{Z}}^{\lambda^{1}-1}\qquad\textnormal{ and }\qquad\widetilde{H}^{n}(F_{f,\mathbf{0}};\ {\mathbb{Z}})=0.$
But there is a great deal more that we can say.
We have
$1=\lambda^{0}=\left(\Gamma^{1}_{f,{\mathbf{z}}}\cdot V\left(\frac{\partial f}{\partial z_{0}}\right)\right)_{\mathbf{0}}\geq\big({\operatorname{mult}}_{\mathbf{0}}\Gamma^{1}_{f,{\mathbf{z}}}\big)\left({\operatorname{mult}}_{\mathbf{0}}\left(\frac{\partial f}{\partial z_{0}}\right)\right),$
by Corollary 12.4 of [5]. Therefore,
${\operatorname{mult}}_{\mathbf{0}}\Gamma^{1}_{f,{\mathbf{z}}}=1,\ \ {\operatorname{mult}}_{\mathbf{0}}\left(\frac{\partial f}{\partial z_{0}}\right)=1,\ \textnormal{ and so }\ {\operatorname{mult}}_{\mathbf{0}}f=2.$
Furthermore, as $\lambda^{0}=1$, the trace $\operatorname{tr}\alpha_{0}$ from Theorem 2.2 must be $\pm 1$, and so ${\operatorname{mult}}_{\mathbf{0}}|\Sigma f|=2$. There are two cases: (1) $\Sigma f$ has a single irreducible component $C$ of multiplicity 2 at the origin or (2) $\Sigma f$ has two smooth irreducible components which are transversely intersected by $V(z_{0})$ at the origin.
In case (1), $\Lambda^{1}=\mu^{\circ}C$ and $\lambda^{1}=({\operatorname{mult}}_{\mathbf{0}}C)\mu^{\circ}_{C}=2\mu^{\circ}_{C}$. But then, by Proposition 2.3, $\tilde{b}_{n-1}=2\mu^{\circ}_{C}-1\leq\mu^{\circ}_{C}$. So we conclude that $\mu^{\circ}_{C}=1$; hence $\lambda^{1}=2$, and $\widetilde{H}^{n-1}(F_{f,\mathbf{0}};\ {\mathbb{Z}})\cong{\mathbb{Z}}$. Furthermore, by Proposition 2.3, the internal monodromy of the vanishing cycles around $C-\{{\mathbf{0}}\}$ must be the identity. ∎
Proof of Proposition 1.4. Suppose $\lambda^{1}=2$ and $\lambda^{0}\geq 2$.
In this case, Theorem 2.1 tells us that, for all primes $\mathfrak{p}$, $\tilde{b}_{n-1}+\,\tau_{\mathfrak{p}}$ equals 0 or $1$; this of course allows for the sum to be 0 for some prime(s) $\mathfrak{p}_{1}$ but equal to $1$ for other prime(s) $\mathfrak{p}_{2}$. There are two cases:
(1) Suppose that $\tilde{b}_{n-1}=0$. Then, for all primes $\mathbf{p}$}, $\tau_{\mathfrak{p}}$ equals 0 or $1$. Thus,
$\widetilde{H}^{n-1}(F_{f,\mathbf{0}};\ {\mathbb{Z}})=0\qquad\textnormal{ and }\qquad\widetilde{H}^{n}(F_{f,\mathbf{0}};\ {\mathbb{Z}})\cong{\mathbb{Z}}^{\lambda^{0}-2}\oplus\big({\mathbb{Z}}/b{\mathbb{Z}}\big)$
for some $b\geq 1$ (here, bb=1 or $b$ is a product of powers of primes $\mathfrak{p}$ such that $\tau_{\mathfrak{p}}=1$). Note that Proposition 1.6 implies $\lambda^{0}\neq 2$.
(2) Suppose that $\tilde{b}_{n-1}\neq 0$. Then $\tilde{b}_{n-1}=1$ and for all primes $\mathfrak{p}$, $\tau_{\mathfrak{p}}=0$. Thus,
$\widetilde{H}^{n-1}(F_{f,\mathbf{0}};\ {\mathbb{Z}})\cong{\mathbb{Z}}\qquad\textnormal{ and }\qquad\widetilde{H}^{n}(F_{f,\mathbf{0}};\ {\mathbb{Z}})\cong{\mathbb{Z}}^{\lambda^{0}-1}.$
Again Proposition 1.6 implies $\lambda^{0}\neq 2$.
However, there are 3 other cases which we can consider, based on the nature of $|\Sigma f|$. (a) $|\Sigma f|$ consists of two smooth irreducible components at the origin, both with $\mu^{\circ}=1$. (b) $|\Sigma f|$ has a single irreducible component $C$ of multiplicity 2 at the origin with $\mu^{\circ}_{C}=1$, and (c) $|\Sigma f|$ has a single smooth component $C$ at the origin such that $\mu^{\circ}_{C}=2$. It is this final case where we have more to say.
In Case (c), $m={\operatorname{mult}}_{\mathbf{0}}|\Sigma f|=1$ and so, by Theorem 2.2, $\operatorname{tr}\alpha_{1}=\pm 1$. However, the characteristic polynomial ${\operatorname{char}}_{\alpha_{1}}(t)$ of $\alpha_{1}$ has degree 2 and is product of cyclotomic polynomials. It follows that ${\operatorname{char}}_{\alpha_{1}}(t)=t^{2}\pm t+1$, which are both irreducible over $\mathbb{Z}$} (or $\mathbb{Q}$}). It follows that we cannot be in Case (2) from above, for if $\widetilde{H}^{n-1}(F_{f,\mathbf{0}};\ {\mathbb{Z}})\cong{\mathbb{Z}}$, then the characteristic polynomial of the Milnor monodromy action on $\widetilde{H}^{n-1}(F_{f,\mathbf{0}};\ {\mathbb{Z}})$ would have to divide ${\operatorname{char}}_{\alpha_{1}}(t)$.
Therefore, if $\lambda^{1}=2$, $\lambda^{0}\geq 2$, and $m=1$, we conclude that we are in Case (2), $\lambda^{0}\geq 3$,
$\widetilde{H}^{n-1}(F_{f,\mathbf{0}};\ {\mathbb{Z}})=0\qquad\textnormal{ and }\qquad\widetilde{H}^{n}(F_{f,\mathbf{0}};\ {\mathbb{Z}})\cong{\mathbb{Z}}^{\lambda^{0}-2}\oplus\big({\mathbb{Z}}/b{\mathbb{Z}}\big)$
for some $b\geq 1$. However, there is more that we say about the prime powers that divide $b$.
For each prime $\mathfrak{p}$, the question is: does ${\operatorname{char}}_{\alpha_{1}}(t)=t^{2}\pm t+1$ factor in $({\mathbb{Z}}/\mathfrak{p}{\mathbb{Z}})[t]$? A little number theory tells one that if ${\operatorname{char}}_{\alpha_{1}}(t)=t^{2}\pm t+1$ factors in $({\mathbb{Z}}/\mathfrak{p}{\mathbb{Z}})[t]$, then $\mathfrak{p}$ is either $3$ or congruent to $1$ modulo $6$. Thus in the isomorphism $\widetilde{H}^{n}(F_{f,\mathbf{0}};\ {\mathbb{Z}})\cong{\mathbb{Z}}^{\lambda^{0}-2}\oplus\big({\mathbb{Z}}/b{\mathbb{Z}}\big)$, either $b=1$ or $b$ is a product of prime powers where the primes are either $3$ or congruent to $1$ modulo $6$. ∎
Proof of Proposition 1.5. Suppose $\lambda^{0}=2$ and $\lambda^{1}\geq 3$.
Unsurprisingly, the analysis here is analogous to the $\lambda^{1}=2$ situation above. In our current case, Theorem 2.1 tells us that all primes $\mathfrak{p}$, $\tilde{b}_{n}+\,\tau_{\mathfrak{p}}$ equals 0 or 1. There are two cases.
(1) Suppose that $\tilde{b}_{n}=0$. Then, for all primes $\mathbf{p}$}, $\tau_{\mathfrak{p}}$ equals 0 or $1$. Thus,
$\widetilde{H}^{n-1}(F_{f,\mathbf{0}};\ {\mathbb{Z}})\cong{\mathbb{Z}}^{\lambda^{1}-2}\qquad\textnormal{ and }\qquad\widetilde{H}^{n}(F_{f,\mathbf{0}};\ {\mathbb{Z}})\cong{\mathbb{Z}}/m{\mathbb{Z}}$
for some $m\geq 1$.
(2) Suppose that $\tilde{b}_{n}\neq 0$. Then $\tilde{b}_{n}=1$ and for all primes $\mathfrak{p}$, $\tau_{\mathfrak{p}}=0$. Thus,
$\widetilde{H}^{n-1}(F_{f,\mathbf{0}};\ {\mathbb{Z}})\cong{\mathbb{Z}}^{\lambda^{1}-1}\qquad\textnormal{ and }\qquad\widetilde{H}^{n}(F_{f,\mathbf{0}};\ {\mathbb{Z}})\cong{\mathbb{Z}}.$
Now assume that $m={\operatorname{mult}}_{\mathbf{0}}\Sigma f=2$. Then, analogous to what we saw in the $\lambda^{1}=2$ situation, Theorem 2.2 implies that $\operatorname{tr}\alpha_{0}=\pm 1$ and so ${\operatorname{char}}_{\alpha_{0}}(t)=t^{2}\pm t+1$. Thus we must be in Case (1) above and either $m=1$ or $m$ is a product of prime powers where the primes are either $3$ or congruent to $1$ modulo $6$. ∎
Proof of Proposition 1.6. Suppose $\lambda^{0}=\lambda^{1}=3$.
Theorem 2.1 tells us that $\tilde{b}_{n-1}<3$. Proposition 1.6 implies $\tilde{b}_{n-1}\neq 0,1$. Thus, we must have $\tilde{b}_{n-1}=\tilde{b}_{n}=2$. But then Theorem 2.1 also tells us that for all primes $\mathfrak{p}$, $\tau_{\mathfrak{p}}=0$, i.e., there can be no torsion. Thus
$\widetilde{H}^{n-1}(F_{f,\mathbf{0}};\ {\mathbb{Z}})\cong{\mathbb{Z}}^{2}\qquad\textnormal{ and }\qquad\widetilde{H}^{n}(F_{f,\mathbf{0}};\ {\mathbb{Z}})\cong{\mathbb{Z}}^{2}.$
Now suppose that $|\Sigma f|$ has exactly one irreducible component $C$. Then ${\operatorname{mult}}_{\mathbf{0}}C\neq 3$, for if it were, then $\mu^{\circ}_{C}$ would have to be 1, and Proposition 2.3 would tell us that $\tilde{b}_{n-1}\leq 1$; a contradiction. It is also true that ${\operatorname{mult}}_{\mathbf{0}}C\neq 2$, for if it were $\lambda^{1}$ would be even; a contradiction. Therefore $C$ is smooth and transversely intersected by $V(z_{0})$ at $\mathbf{0}$}, and $\mu^{\circ}_{C}$ must be 3. ∎
4. Examples or the lack thereof
Here, we wish to give examples of the various cases from the introduction or say that no such examples are known (to us).
Example 4.1.
Consider the case where $\lambda^{0}=0$. Examples of this are trivial to produce, simply take a product of an isolated singularity with a complex line, e.g., $f(z_{0},z_{1},z_{2})=z_{1}^{2}+z_{2}^{2}$.
Example 4.2.
Consider the case where $\lambda^{1}=1$ and $\lambda^{0}\geq 1$. It is easy to produce examples of this case, for instance $f=z_{2}^{2}-z_{1}^{3}-z_{0}z_{1}^{2}$.
Remark 4.3.
Consider the case where $\lambda^{0}=1$ and $\lambda^{1}\geq 2$. We do not know if this case can actually occur. Certainly it cannot occur for a smooth critical locus, since we showed $m={\operatorname{mult}}_{\mathbf{0}}|\Sigma f|$ must be 2.
Example 4.4.
Consider the case where $\lambda^{1}=2$ and $\lambda^{0}\geq 2$. We showed that $\lambda^{0}$ must be at least 3 in this case. It is easy to produce examples in all 3 subcases (a), (b), and (c) from the proof of Proposition 1.4. We have the examples, in order,
$f=(z_{0}^{2}-z_{1}^{2}+z_{2}^{2})z_{2},\qquad g=(z_{1}^{2}-z_{0}^{3}+z_{2})z_{2},\ \ \textnormal{ and }\quad h=z_{2}^{2}-z_{1}^{4}-z_{0}z_{1}^{3}.$
Note that, as $h$ is the “suspension” of $z_{1}^{4}-z_{0}z_{1}^{3}$, there is no torsion in the cohomology of the Milnor fiber of $h$ even though Proposition 1.4 allows for it to exist.
Example 4.5.
Consider the case where $\lambda^{0}=2$ and $\lambda^{1}\geq 3$. Examples of this are not so easy to produce, but they exist. For ease, let us denote our coordinates by $(x,y,z)$ in place of $(z_{0},z_{1},z_{2})$.
Let $f=(z^{2}-x^{2}-y^{2})(z-x)$. Then, we have $\Sigma f=V(y,z-x)$, and
$V\left(\frac{\partial f}{\partial z}\right)=V(2z(z-x)+(z^{2}-x^{2}-y^{2})).$
We find, as cycles, that
$V\left(\frac{\partial f}{\partial y},\frac{\partial f}{\partial z}\right)=V(y,3z+x)+3V(y,z-x),$
so that $\Gamma^{1}_{f,{\mathbf{z}}}=V(y,3z+x)$ and $\Lambda^{1}_{f,{\mathbf{z}}}=3V(y,z-x)$. It is now easy to show that $\lambda^{0}=2$ and $\lambda^{1}=3$.
In fact, we can show that we are in Case (1) of Proposition 1.5. Up to analytic isomorphism, $f$ is the homogeneous polynomial $f=(zx-y^{2})z$. Consequently, $F_{f,{\mathbf{0}}}$ is diffeomorphic to $f^{-1}(1)$, i.e., to the space where $(zx-y^{2})z=1$; thus $z\neq 0$ and $x=(y^{2}+1/z)/z$. Thus, $F_{f,{\mathbf{0}}}$ is homotopy equivalent to $S^{1}$.
Remark 4.6.
Finally, consider the case where $\lambda^{0}=\lambda^{1}=3$. Once again, we do not know if this case can actually occur.
5. Remarks
Remark 5.1.
Given the result of Siersma in [12], we suspected that Theorem 2.1 should be the correct generalization on the level of cohomology. Our original attack on proving such a generalization was to show that a non-empty relative polar curve necessarily implies that the internal monodromy in Proposition 2.3 cannot be the identity. However, the example/counterexample of $f=z_{2}^{2}-z_{1}^{3}-z_{0}^{2}z_{1}^{2}$ shows that this is not true.
Remark 5.2.
At the end of Example 5.1 of [11], we address the case where $\lambda^{1}=2$, $\lambda^{0}\neq 0$, and ${\operatorname{mult}}_{\mathbf{0}}|\Sigma f|=1$. There we concluded quickly that there can be no torsion in the cohomology of the Milnor fiber. While it is possible that this is true, certainly we did not give a careful argument that proved it. As of now, the best we can say is what we wrote in Proposition 1.4.
Remark 5.3.
Throughout this article, we have allowed for torsion in the cohomology of the Milnor fiber. In fact, we do not know if such torsion can possibly exist when $\dim_{\mathbf{0}}\Sigma f=1$.
However, when $\dim_{\mathbf{0}}\Sigma f=2$, there are the famous examples of Cohen, Denham, and Suciu in [3] of hypersurfaces with torsion in the cohomology of the Milnor fiber.